# Cornwall-Jackiw-Tomboulis effective field theory to nonuniversal equation of state of an ultracold Bose gas

### v1

@@ IV Frequency shifts induced by NONUNIVERSAL beyond-LHY EOS
In Sec. III, we have rigorously derived the nonuniversal next-to-LHY terms in EOS (see Eqs. (14) and (15)) to the order of $\left(\rho a_{\text{s}}^{3}\right)^{2}$. The purpose of Sec. IV is to propose an experimental protocol to observe the nonuniversal beyond-LHY corrections to the EOS by calculating the frequency shifts in the breathing modes. In contemporary experiments, the gas parameter $\rho a_{\text{s}}^{3}$ typically remains below $10^{-4}$, rendering even the initial LHY correction to the mean-field energy, scaling as $\sqrt{\rho a_{\text{s}}^{3}}$, negligible at best, contributing less than $1\%$ of the total energy. Consequently, the higher-order corrections beyond LHY in Eq. (15) are anticipated to be too subtle to discern in density profiles or release energies. While enhancing quantum fluctuations by tuning $a_{\text{s}}$ via Feshbach resonance [31, 32, 50, 33, 35, 34] can amplify their impact, we propose an alternative route: studying the frequency shifts in collective excitations [51, 52, 53, 54] induced by quantum fluctuations [55]. To this end, we augment our model system with a 3D harmonic trap, defined by $V_{\text{ext}}=\frac{1}{2}m\omega^{2}_{0}r^{2}$, and embark on calculating the resulting shifts in the breathing mode frequency, attributed to beyond-LHY terms in Eq. (15). This endeavor is motivated by the potential to experimentally verify and quantify the nonuniversal contributions to the EOS.
@@
As an initial step, utilizing Eq. (15), we can iteratively derive the expansion of the ground state, expressed in terms of $\alpha=\left(32/3\sqrt{\pi}\right)a_{\text{s}}^{3/2}$,
@@
Equation (20) simplifies into the basic hydrodynamic equation $m\omega^{2}\delta\rho+\nabla\cdot\left[\rho_{\text{TF}}\nabla g_{0}\delta\rho\right]=0$ in case of $\alpha=0$. For this simplified scenario, the calculated frequency $\omega$ exhibits analytical solutions of the form $\omega\left(n_{r},l\right)=\omega_{0}\left(2n_{r}^{2}+2n_{r}l+3n_{r}+l\right)^{1/2}$, with $n_{r}$ representing the number of radial nodes and $l$ denoting the angular momentum associated with the excitation. This analytical expression provides a direct link between the frequency of the excitation and its quantum numbers.
@@
Equation (22) constitutes another pivotal finding, encapsulating corrections to the breathing mode frequency that scale with the gas parameter $\rho\left(0\right)a_{\text{s}}^{3}$. The corrections in Eq. (22) stem from diverse origins, encompassing both LHY and beyond-LHY contributions to the EOS. The first term on the right-hand side of Eq. (22) represents the ubiquitous LHY-mediated fractional shift of the breathing mode frequency, which scales as $\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{1/2}$. All the subsequent terms in Eq. (22) are of higher order than $\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{1/2}$, offering insights into more intricate effects. Notably, among these beyond-LHY terms, the one scaling with $\left[\rho\left(0\right)a_{\text{s}}^{3}\right]$ is a so-called next-to-LHY universal fractional shift. This term, specifically $\frac{-16384}{135\pi}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]$, underscores the persistence of universal behavior beyond the leading LHY correction. Particularly, terms in Eq. (22) involving $r_{\text{s}}$ represent nonuniversal contributions, arising from the finite-range effective potential, further illuminating the intricacies of the system’s response. Thus, the complexity deepens with the term $\frac{3682-\frac{2205}{16}\pi^{2}\frac{r_{\text{s}}}{a_{\text{s}}}}{3\sqrt{\pi}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{3/2}$ in Eq. (22), which encapsulates both the next-next-to-LHY universal fractional shift $\frac{3682}{3\sqrt{\pi}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{3/2}$ and a next-to-LHY nonuniversal contribution $\frac{-\frac{2205}{16}\pi^{2}\frac{r_{\text{s}}}{a_{\text{s}}}}{3\sqrt{\pi}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{3/2}$. Similarly, the last term $\left(\frac{-71200}{\pi^{2}}+6465\frac{r_{\text{s}}}{a_{\text{s}}}\right)\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{2}$ in Eq. (22) comprises two distinct components: the next-next-next-to-LHY universal fractional shift $\frac{-71200}{\pi^{2}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{2}$ and the next-next-to-LHY nonuniversal fractional shift $6465\frac{r_{\text{s}}}{a_{\text{s}}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{2}$. We remark that to our best knowledge, the analytical expressions of beyond-LHY-induced fractional shift of the breathing mode frequency in Eq. (22) are obtained for the first time.
@@
We are now poised to discuss the influence of the nonuniversal EOS in Eq. (15), stemming from the finite-range interaction parameter $r_{\text{s}}$, on the fractional shift of the breathing mode frequency in Eq. (22). To visualize this effect, we have plotted frequency shifts of the breathing mode $\delta\omega/\omega$ in Eq. (22) into Fig. 1, showing how the dimensionless finite-range interaction parameter of $r_{\text{s}}/a_{\text{s}}$ can affect the frequency shifts of $\delta\omega/\omega$. Note that the red solid curve in Fig. 1 corresponds to case of vanishing the finite-range interaction by taking the dimensionless parameter of $r_{\text{s}}/a_{s}=0$; meanwhile the black dot-dashed and blue dashed curves in Fig. 1 demonstrate how the repulsive ($r_{\text{s}}/a_{\text{s}}=0.75$) and attractive ($r_{\text{s}}/a_{\text{s}}=-0.75$) finite-range interaction can affect frequency shifts of $\delta\omega/\omega$. As shown in Fig. 1, the finite-range interaction has relative small effects on frequency shifts of $\delta\omega/\omega$ compared with the case of vanishing the finite-range interaction when the gas parameter of $\rho a^{3}_{\text{s}}$ is small. In contrast, with the increase of the $\rho a^{3}_{\text{s}}$, the effect of finite-range interaction on frequency shifts of $\delta\omega/\omega$ becomes to be significant. For example, at the typical experimental parameter onset of $\rho a^{3}_{\text{s}}\sim 10^{-3}$, the derivation of the fractional shift of $\delta\omega/\omega$ in case of $r_{\text{s}}/a_{\text{s}}=0.75$ from the case of $r_{\text{s}}/a_{\text{s}}=0$ is calculated as more than $0.5\%$, showing the finite-range effect well in reach in experiments. We point out that a precision of $<0.3\%$ in measuring collective frequencies has already been established [19, 20, 21, 22, 23], offering opportunities to probe the nonuniversial beyond-LHY corrections to EOS experimentally.
@@ V CONCLUSION and outlook
In summary, we have theoretically investigated the nonuniversal EOS for a weakly-interacting Bose gas with the finite-range interatomic interaction. With the framework of CJT effective field theory under the two-loop approximation, we obtain analytical expressions for quantum depletion and chemical potential of model system, representing the next-to-LHY correction to nonuniversal EOS induced by finite-range effects. These analytical results represent significant generalizations of the nonuniversal LHY EOS studied in Refs. [43, 26, 30] to the beyond LHY regimes, offering fresh insights into understanding the quantum behavior induced by the quantum fluctuations in many-body bosonic systems. We further calculate the frequency shifts in the breathing mode induced by the nonuniversal beyond-LHY EOS. Therefore, the beyond-LHY effects studied in this work should be observable within the current experiment capability.
@@ Appendix A CJT EFFECTIVE FIELD THEORY
$\mathcal{Z}\left[J,K\right]$ $=e^{-W\left[J,K\right]/\hbar}=\int\mathcal{D}\left[\Psi\right]\exp\left\{-\frac{1}{\hbar}\left[\int d^{4}x\mathcal{L}\left[\Psi\right]+J\left(x\right)\Psi\left(x\right)+\frac{1}{2}\int d^{4}xd^{4}y\Psi\left(x\right)K\left(x,y\right)\Psi\left(y\right)\right]\right\}.$ (23)
@@
$\frac{\delta W\left[J,K\right]}{\delta J\left(x\right)}$ $=$ $-\frac{\delta\ln\mathcal{Z}}{\delta J\left(x\right)}=\left\langle\Psi\left(x\right)\right\rangle_{J,K}=\phi\left(x\right),$ (24a)
$\frac{\delta W\left[J,K\right]}{\delta K\left(x,y\right)}$ $=$ $-\frac{\delta\ln\mathcal{Z}}{\delta K\left(x,y\right)}=\frac{1}{2}\left\langle\Psi\left(x\right)\Psi\left(y\right)\right\rangle_{J,K}$ (24b)
@@
$\mathcal{Z}$ $\simeq e^{-S\left[\phi,J,K\right]/\hbar}\int\mathcal{D}[\widetilde{\Psi}]e^{-\frac{1}{2\hbar}\int d^{4}xd^{4}y\tilde{\Psi}\left(x\right)\left[D_{0}^{-1}+K\right]\tilde{\Psi}\left(y\right)}$
@@
To ensure the self-consistency and comprehensiveness of this work, we provide a concise overview of the crucial steps involved in deriving the self-consistent $V_{\text{eff}}$ that fulfills the condition of gapless excitations (see e.g. Refs. [58, 18] and the references therein). In particular, we delve into the derivation of the effective potential $V_{\text{eff}}$ as presented in Eq. (8). To effectively implement the CJT effective field theory [39], it is advantageous to initially transform the CJT effective potential density into the momentum-frequency space
$\phi_{1}$$\left(\mathbf{r},\tau\right)$ $=$ $\sqrt{\frac{1}{\beta\hbar\mathcal{V}}}\sum_{{\mathbf{k}},\omega_{n}}e^{i{\mathbf{k}}\cdot{\mathbf{r}}+i\omega_{n}\tau/\hbar}\phi_{1}\left({\mathbf{k}},\omega_{n}\right),$ (32a)
$\phi_{2}$$\left(\mathbf{r},\tau\right)$ $=$ $\sqrt{\frac{1}{\beta\hbar\mathcal{V}}}\sum_{{\mathbf{k}},\omega_{n}}e^{i{\mathbf{k}}\cdot{\mathbf{r}}+i\omega_{n}\tau/\hbar}\phi_{2}\left({\mathbf{k}},\omega_{n}\right),$ (32b)
@@
Based on the CJT effective field theory, we find the CJT effective potential density from the Lagrangian density (2) as
@@ B.1 Zero-loop
In zero-loop approximation, we solely consider the first diagram depicted in Fig. 2 as the contribution to $\Phi\left[\phi_{0}\right]$. Consequently, the effective potential can be succinctly expressed as $V_{\text{eff}}[\phi_{0}]=-\frac{\mu}{2}\phi_{0}^{2}+\frac{g_{0}}{8}\phi_{0}^{4}$. Notably, $V_{\text{eff}}[\phi_{0}]$ involves a single variable $\phi_{0}$, which represents the mean-field. By minimizing the $V_{\text{eff}}[\phi_{0}]$ with respect to $\phi_{0}$ and incorporating thermodynamic relationships, we derive $\mu=g_{0}\rho$ and $\rho=\frac{g_{0}}{2}\phi_{0}^{2}$. These results align perfectly with the well-established findings within the mean-field approximation framework.
@@ B.2 One-loop
being the inversion propagator within one-loop approximation. Minimizing the effective potential density $V_{\text{eff}}\left[\phi_{0},G\right]$ with respect to the elements of the propagator $G\left(k\right)$, we obtain that $G^{-1}\left(k\right)=G_{0}^{-1}\left(k\right)$. Consequently, we reformulate $V_{\text{eff}}[\phi_{0},G]$ as $V_{\text{eff}}[\phi_{0},G_{0}]$, which in this approximation takes the form: $V_{\text{eff}}[\phi_{0},G_{0}]=-\frac{\mu}{2}\phi_{0}^{2}+\frac{g_{0}}{8}\phi_{0}^{4}+\frac{1}{2}\int_{\beta}\text{\text{Tr}}\left[\ln G^{-1}_{0}\left(k\right)\right]$. Further minimization of $V_{\text{eff}}[\phi_{0},G_{0}]$ with respect to order parameter $\phi_{0}$ and application of thermodynamic relationships lead to the quantum depletion: $\rho_{\text{ex}}=\frac{8\rho}{3\sqrt{\pi}}\left(\rho a_{\text{s}}^{3}\right)^{1/2}-64\sqrt{\pi}\rho\frac{r_{\text{s}}}{a_{\text{s}}}\left(\rho a_{\text{s}}^{3}\right)^{3/2}$. This result aligns with the findings reported in Ref. [30], thereby validating the one-loop approximation approach.

### v2

@@ IV Frequency shifts induced by NONUNIVERSAL beyond-LHY EOS
In Sec. III, we have rigorously derived the nonuniversal next-to-LHY term in EOS (see Eqs. (14) and (15)) to the order of $\left(\rho a_{\text{s}}^{3}\right)^{2}$. The purpose of Sec. IV is to propose an experimental protocol to observe the nonuniversal beyond-LHY corrections to the EOS by calculating the frequency shifts in the breathing modes. In contemporary experiments, the gas parameter $\rho a_{\text{s}}^{3}$ typically remains below $10^{-4}$, rendering even the initial LHY correction to the mean-field energy, scaling as $\sqrt{\rho a_{\text{s}}^{3}}$, negligible at best, contributing less than $1\%$ of the total energy. Consequently, the higher-order corrections beyond LHY in Eq. (15) are anticipated to be too subtle to discern in density profiles or release energies. While enhancing quantum fluctuations by tuning $a_{\text{s}}$ via Feshbach resonance [31, 32, 55, 33, 35, 34] can amplify their impact, we propose an alternative route: studying the frequency shifts in collective excitations [56, 57, 58, 59] induced by quantum fluctuations [60]. To this end, we augment our model system with a 3D harmonic trap, defined by $V_{\text{ext}}=\frac{1}{2}m\omega^{2}_{0}r^{2}$, and embark on calculating the resulting shifts in the breathing mode frequency, attributed to beyond-LHY terms in Eq. (15). This endeavor is motivated by the potential to experimentally verify and quantify the nonuniversal contributions to the EOS.
@@
As an initial step, utilizing Eq. (15), we can iteratively derive the expansion of the ground state density, expressed in terms of $\alpha=\left(32/3\sqrt{\pi}\right)a_{\text{s}}^{3/2}$,
@@
Equation (20) simplifies into the basic hydrodynamic equation $m\omega^{2}\delta\rho+\nabla\cdot\left[g_{0}\rho_{\text{TF}}\nabla\delta\rho\right]=0$ in case of $\alpha=0$. For this simplified scenario, the calculated frequency $\omega$ exhibits analytical solutions of the form $\omega\left(n_{r},l\right)=\omega_{0}\left(2n_{r}^{2}+2n_{r}l+3n_{r}+l\right)^{1/2}$, with $n_{r}$ representing the number of radial nodes and $l$ denoting the angular momentum associated with the excitation. This analytical expression provides a direct link between the frequency of the excitation and its quantum numbers.
@@
Equation (22) constitutes another pivotal finding, encapsulating corrections to the breathing mode frequency that scale with the gas parameter $\rho\left(0\right)a_{\text{s}}^{3}$. The corrections in Eq. (22) stem from diverse origins, encompassing both LHY and beyond-LHY contributions to the EOS. The first term on the right-hand side of Eq. (22) represents the ubiquitous LHY-mediated fractional shift of the breathing mode frequency, which scales as $\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{1/2}$. All the subsequent terms in Eq. (22) are of higher order than $\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{1/2}$, offering insights into more intricate effects. Notably, among these beyond-LHY terms, the one scaling with $\left[\rho\left(0\right)a_{\text{s}}^{3}\right]$ is a so-called next-to-LHY universal fractional shift. This term, specifically $\frac{-16384}{135\pi}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]$, underscores the persistence of universal behavior beyond the leading LHY correction. Particularly, terms in Eq. (22) involving $r_{\text{s}}$ represent nonuniversal contributions, arising from the finite-range effective potential, further illuminating the intricacies of the system’s response. Thus, the complexity deepens with the term $\frac{3682-\frac{2205}{16}\pi^{2}\frac{r_{\text{s}}}{a_{\text{s}}}}{3\sqrt{\pi}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{3/2}$ in Eq. (22), which encapsulates both the next-next-to-LHY universal fractional shift $\frac{3682}{3\sqrt{\pi}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{3/2}$ and a LHY nonuniversal contribution $\frac{-\frac{2205}{16}\pi^{2}\frac{r_{\text{s}}}{a_{\text{s}}}}{3\sqrt{\pi}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{3/2}$. Similarly, the last term $\left(\frac{-71200}{\pi^{2}}+6465\frac{r_{\text{s}}}{a_{\text{s}}}\right)\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{2}$ in Eq. (22) comprises two distinct components: the next-next-next-to-LHY universal fractional shift $\frac{-71200}{\pi^{2}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{2}$ and the next-to-LHY nonuniversal fractional shift $6465\frac{r_{\text{s}}}{a_{\text{s}}}\left[\rho\left(0\right)a_{\text{s}}^{3}\right]^{2}$. We remark that to our best knowledge, the analytical expressions of beyond-LHY-induced fractional shift of the breathing mode frequency in Eq. (22) are obtained for the first time.
@@
We are now poised to discuss the influence of the nonuniversal EOS in Eq. (15), stemming from the finite-range interaction parameter $r_{\text{s}}$, on the fractional shift of the breathing mode frequency in Eq. (22). To visualize this effect, we have plotted frequency shifts of the breathing mode $\delta\omega/\omega$ in Eq. (22) into Fig. 1, showing how the dimensionless finite-range interaction parameter of $r_{\text{s}}/a_{\text{s}}$ can affect the frequency shifts of $\delta\omega/\omega$. Note that the red solid curve in Fig. 1 corresponds to case of vanishing the finite-range interaction by taking the dimensionless parameter of $r_{\text{s}}/a_{s}=0$; meanwhile the black dot-dashed and blue dashed curves in Fig. 1 demonstrate how the attractive ($r_{\text{s}}/a_{\text{s}}=0.75$) and repulsive ($r_{\text{s}}/a_{\text{s}}=-0.75$) finite-range interaction can affect frequency shifts of $\delta\omega/\omega$. As shown in Fig. 1, the finite-range interaction has relative small effects on frequency shifts of $\delta\omega/\omega$ compared with the case of vanishing the finite-range interaction when the gas parameter of $\rho a^{3}_{\text{s}}$ is small. In contrast, with the increase of the $\rho a^{3}_{\text{s}}$, the effect of finite-range interaction on frequency shifts of $\delta\omega/\omega$ becomes to be significant. For example, at the typical experimental parameter onset of $\rho a^{3}_{\text{s}}\sim 10^{-3}$, the derivation of the fractional shift of $\delta\omega/\omega$ in case of $r_{\text{s}}/a_{\text{s}}=0.75$ from the case of $r_{\text{s}}/a_{\text{s}}=-0.75$ is calculated as more than $0.5\%$, showing the finite-range effect well in reach in experiments. We point out that a precision of $<0.3\%$ in measuring collective frequencies has already been established [19, 20, 21, 22, 23], offering opportunities to probe the nonuniversial beyond-LHY corrections to EOS experimentally.
@@ V CONCLUSION and outlook
In summary, we have theoretically investigated the nonuniversal EOS for a weakly-interacting Bose gas with the finite-range interatomic interaction. With the framework of CJT effective field theory under the two-loop approximation, we obtain analytical expressions for quantum depletion and chemical potential of model system, representing the next-to-LHY correction to nonuniversal EOS induced by finite-range effects. These analytical results represent significant generalizations of the nonuniversal LHY EOS studied in Refs. [43, 26, 30] to the beyond-LHY regimes, offering fresh insights into understanding the quantum behavior induced by the quantum fluctuations in many-body bosonic systems. We further calculate the frequency shifts in the breathing mode induced by the nonuniversal beyond-LHY EOS. Therefore, the beyond-LHY effects studied in this work should be observable within the current experiment capability.
@@ Appendix A CJT EFFECTIVE FIELD THEORY
$\mathcal{Z}\left[J,K\right]$ $=e^{-W\left[J,K\right]/\hbar}=\int\mathcal{D}\left[\mathbf{\Phi}\right]\exp\left\{-\frac{1}{\hbar}\left[\int d^{4}x\mathcal{L}\left[\mathbf{\Phi}\right]+J\left(x\right)\mathbf{\Phi}\left(x\right)+\frac{1}{2}\int d^{4}xd^{4}y\mathbf{\Phi}\left(x\right)K\left(x,y\right)\mathbf{\Phi}\left(y\right)\right]\right\}.$ (23)
@@
$\frac{\delta W\left[J,K\right]}{\delta J\left(x\right)}$ $=$ $-\frac{\delta\ln\mathcal{Z}}{\delta J\left(x\right)}=\left\langle\mathbf{\Phi}\left(x\right)\right\rangle_{J,K}=\phi\left(x\right),$ (24a)
$\frac{\delta W\left[J,K\right]}{\delta K\left(x,y\right)}$ $=$ $-\frac{\delta\ln\mathcal{Z}}{\delta K\left(x,y\right)}=\frac{1}{2}\left\langle\mathbf{\Phi}\left(x\right)\mathbf{\Phi}\left(y\right)\right\rangle_{J,K}$ (24b)
@@
$\mathcal{Z}$ $\simeq e^{-S\left[\phi,J,K\right]/\hbar}\int\mathcal{D}[\widetilde{\mathbf{\Phi}}]e^{-\frac{1}{2\hbar}\int d^{4}xd^{4}y\tilde{\mathbf{\Phi}}\left(x\right)\left[D_{0}^{-1}+K\right]\tilde{\mathbf{\Phi}}\left(y\right)}$
@@
To ensure the self-consistency and comprehensiveness of this work, we provide a concise overview of the crucial steps involved in deriving the self-consistent $V_{\text{eff}}$ that fulfills the condition of gapless excitations (see e.g. Refs. [63, 18] and the references therein). In particular, we delve into the derivation of the effective potential $V_{\text{eff}}$ as presented in Eq. (8). To effectively implement the CJT effective field theory [39], it is advantageous to initially transform the CJT effective potential into the momentum-frequency space
$\phi_{1}$$\left(\mathbf{r},\tau\right)$ $=$ $\sqrt{\frac{1}{\beta\hbar\mathcal{V}}}\sum_{{\mathbf{k}},\omega_{n}}e^{i{\mathbf{k}}\cdot{\mathbf{r}}-i\omega_{n}\tau/\hbar}\phi_{1}\left({\mathbf{k}},\omega_{n}\right),$ (32a)
$\phi_{2}$$\left(\mathbf{r},\tau\right)$ $=$ $\sqrt{\frac{1}{\beta\hbar\mathcal{V}}}\sum_{{\mathbf{k}},\omega_{n}}e^{i{\mathbf{k}}\cdot{\mathbf{r}}-i\omega_{n}\tau/\hbar}\phi_{2}\left({\mathbf{k}},\omega_{n}\right),$ (32b)
@@
Based on the CJT effective field theory, we find the CJT effective potential from the Lagrangian density (2) as
@@ B.1 Zero-loop
In zero-loop approximation, we solely consider the first diagram depicted in Fig. 2 as the contribution to $\Phi\left[\phi_{0}\right]$. Consequently, the effective potential can be succinctly expressed as $V_{\text{eff}}[\phi_{0}]=-\frac{\mu}{2}\phi_{0}^{2}+\frac{g_{0}}{8}\phi_{0}^{4}$. Notably, $V_{\text{eff}}[\phi_{0}]$ involves a single variable $\phi_{0}$, which represents the mean-field. By minimizing the $V_{\text{eff}}[\phi_{0}]$ with respect to $\phi_{0}$ and incorporating thermodynamic relationships, we derive $\mu=g_{0}\rho$ and $\rho_{\text{ex}}=0$. These results align perfectly with the well-established findings within the mean-field approximation framework.
@@ B.2 One-loop
being the inversion propagator within one-loop approximation. Minimizing the effective potential $V_{\text{eff}}\left[\phi_{0},G\right]$ with respect to the elements of the propagator $G\left(k\right)$, we obtain that $G^{-1}\left(k\right)=G_{0}^{-1}\left(k\right)$. Consequently, we reformulate $V_{\text{eff}}[\phi_{0},G]$ as $V_{\text{eff}}[\phi_{0},G_{0}]$, which in this approximation takes the form: $V_{\text{eff}}[\phi_{0},G_{0}]=-\frac{\mu}{2}\phi_{0}^{2}+\frac{g_{0}}{8}\phi_{0}^{4}+\frac{1}{2}\int_{\beta}\text{\text{Tr}}\left[\ln G^{-1}_{0}\left(k\right)\right]$. Further minimization of $V_{\text{eff}}[\phi_{0},G_{0}]$ with respect to order parameter $\phi_{0}$ and application of thermodynamic relationships lead to the quantum depletion: $\rho_{\text{ex}}=\frac{8\rho}{3\sqrt{\pi}}\left(\rho a_{\text{s}}^{3}\right)^{1/2}-64\sqrt{\pi}\rho\frac{r_{\text{s}}}{a_{\text{s}}}\left(\rho a_{\text{s}}^{3}\right)^{3/2}$. This result aligns with the findings reported in Ref. [30], thereby validating the one-loop approximation approach.